# Features are fate: a theory of transfer learning in high-dimensional regression

### v1

@@ Abstract
With the emergence of large-scale pre-trained neural networks, methods to adapt such “foundation” models to data-limited downstream tasks have become a necessity. Fine-tuning, preference optimization, and transfer learning have all been successfully employed for these purposes when the target task closely resembles the source task, but a precise theoretical understanding of “task similarity” is still lacking. While conventional wisdom suggests that simple measures of similarity between source and target distributions, such as $\phi$-divergences or integral probability metrics, can directly predict the success of transfer, we prove the surprising fact that, in general, this is not the case. We adopt, instead, a feature-centric viewpoint on transfer learning and establish a number of theoretical results that demonstrate that when the target task is well represented by the feature space of the pre-trained model, transfer learning outperforms training from scratch. We study deep linear networks as a minimal model of transfer learning in which we can analytically characterize the transferability phase diagram as a function of the target dataset size and the feature space overlap. For this model, we establish rigorously that when the feature space overlap between the source and target tasks is sufficiently strong, both linear transfer and fine-tuning improve performance, especially in the low data limit. These results build on an emerging understanding of feature learning dynamics in deep linear networks, and we demonstrate numerically that the rigorous results we derive for the linear case also apply to nonlinear networks.
@@ 1 Introduction
State of the art neural network models have billions to trillions of parameters and are trained on datasets of a similar scale. The benefits of dataset scale are manifest in the astounding generalization capability of these foundation models (Bahri et al., 2024). For many applications, however, datasets of the scale used for natural language processing or computer vision are hard, if not impossible, to generate.
To alleviate the problem of inadequate dataset scale, the representations of a foundation model seem to provide a useful inductive bias for adaptation to a target task.
@@
While they are now ubiquitous, transfer learning methods lack a solid theoretical foundation or algorithmic design principles. As such, it remains difficult to predict when—and with which approach—transfer learning will outperform training on the target task alone. Intuitively, if the source task resembles the target task, transfer learning should be beneficial. The important question of how to quantify task relatedness is one that remains unanswered. In this work, we address this question and prove the surprising fact that discrepancies between source and target data distributions can be misleading when it comes to transferability. We instead find that the feature space learned during pretraining is the relevant object for predicting transfer performance, which means that model-agnostic metrics between tasks are unlikely to successfully predict task overlap. Of course, adopting a feature-centric viewpoint creates model-specific challenges because unambiguously identifying learned features remains an outstanding and difficult characterization problem for deep neural networks . For this reason, in this work we focus on deep linear networks trained with gradient flow, as feature learning dynamics are well-understood in this setting. We develop an intuitive understanding of linear transfer and full fine-tuning in this model. In contrast to other recent work, we quantify transfer performance relative to training on the target task alone and precisely identify when transfer learning leads to improved performance, effectively building a phase diagram for transfer efficiency. Finally, we show in numerical experiments that this picture holds qualitatively for nonlinear networks, as well.
1.1 Related work
@@
  • Theoretical aspects of transfer learning A number of recent works have studied theoretical aspects of transfer learning, focusing on the risk associated with various transfer algorithms. Wu et al. (2020) use information theory to derive bounds on the risk of transfer learning using a mixture of source and target data. Shilton et al. (2017) analyze transfer in the context of gaussian process regression. Tripuraneni et al. (2020) work in a fairly general setting, and derive bounds on the generalization error of transferred models through a complexity argument, highlighting the importance of feature diversity among tasks. Liu et al. (2019); Neyshabur et al. (2020) study transfer learning from the perspective of the loss landscape and find that transferred models often find flatter minima than those trained from scratch. Consistent with our feature-centric viewpoint, Kumar et al. (2022) show that fine-tuning can distort the pretrained features, leading to poor out of distribution behavior.
@@
  • Transfer learning in solvable models Similar to our approach, several theory works have worked with analytically tractable models to more precisely characterize transfer performance. Lampinen & Ganguli (2018); Atanasov et al. (2021); Shachaf et al. (2021) also study transfer learning in deep linear networks, but focus on the generalization error alone, not the transferability relative to a scratch trained baseline, which obfuscates the conditions for transfer learning to be beneficial. Gerace et al. (2022) studies transfer learning with small nonlinear networks with data generated from a “hidden manifold“ (Goldt et al., 2020) and find transfer learning to be effective when tasks are very similar, and data is scarce, but do not theoretically describe regions of negative transfer. Saglietti & Zdeborova (2022) studies knowledge distillation in a solvable model, which can be viewed as a special case of transfer learning.
  • Metrics for transferability In an effort to define task similarity, there has been recent interest (Gao & Chaudhari, 2021; Nguyen et al., 2021; Alvarez-Melis & Fusi, 2020) in designing metrics to predict transfer performance, an approach whose validity we assess in Section 2.1.
  • Feature learning The notion of feature learning is central to our results. While the rich, feature learning regime is often heuristically defined as the opposite of the neural tangent regime (Jacot et al., 2018), a precise definition is still lacking. Nevertheless, there has been an explosion of interest in understanding dynamics in these two regimes of neural network optimization Woodworth et al. (2020); Atanasov et al. (2021); Yang & Hu (2021); Kunin et al. (2024); Yun et al. (2021); Chizat (2020) focus on the role of initialization, learning rate, and implicit bias in feature learning. Petrini et al. (2022) highlights the potential for overfitting when training in the feature learning regime.
@@ 2 General theoretical setting
We begin by introducing the general theoretical framework under which we study transfer learning. We consider source and target regression tasks defined by probability distributions $p_{\mathrm{s}}(\bm{x},y)$ and $p_{\mathrm{t}}(\bm{x},y)$ over inputs $\bm{x}\in\mathbb{R}^{d}$ and labels $y\in\mathbb{R}$. We focus on label shift, in which $p_{\mathrm{s}}(\bm{x},y)=p(\bm{x})p_{\mathrm{s}}(y\mid\bm{x})$ and $p_{\mathrm{t}}(\bm{x},y)=p(\bm{x})p_{\mathrm{t}}(y\mid\bm{x})$ for the same input distribution $p(\bm{x})$. The labels are generated from noisy samples of source and target functions
$y_{\textrm{s}}$ $=f_{\textrm{s}}^{*}(\bm{x})+\epsilon$ (1)
$y_{\textrm{t}}$ $=f_{\textrm{t}}^{*}(\bm{x})+\epsilon$ (2)
where $f_{\textrm{s}}^{*}(\bm{x}),f_{\textrm{t}}^{*}(\bm{x})\in L_{2}(p(\bm{x}))$ and $\epsilon\sim\mathcal{N}(0,\sigma^{2})$. During pretraining, we train a model with parameters $\bm{\Theta}=(\bm{c},\bm{\theta})$ of the form
@@
where $\hat{\mathbb{E}}_{p}(h(\bm{x},y))=\frac{1}{n}\sum_{i=1}^{n}h(\bm{x}_{i},y_{i})$ is the expectation over the empirical distribution of $\mathcal{D}$. We focus on two widely employed transfer methods, linear transfer and fine-tuning. In linear transfer, the pretrained features $\phi(\bm{x},\bm{\theta})$ are frozen and only the output weights $\bm{c}$ are trained on the target task. In fine-tuning, the entire set of parameters $\bm{\Theta}$ are trained from the pretrained initialization on the target task. To optimize the loss functions (4) and (5), we use gradient flow,
@@
where we have set the learning rate equal to unity for the purpose of analysis. To assess the performance of transfer learning we compare the performance of the transferred model to a scratch-trained model with the same architecture (3) trained only on the target task from a random initialization. We introduce the transferability to quantify this relationship:
@@
$\mathcal{R}=\mathbb{E}_{p(\bm{x})}[(f(\bm{x},\bm{\Theta})-f^{*}(\bm{x})^{2}].$ (8)
@@
2.1 Dataset similarity is not predictive of transfer efficiency
The common wisdom in transfer learning is that related tasks should transfer effectively to one another. However, a standard and mathematically precise definition of task relatedness is currently lacking. One reasonable notion of task relatedness is to compare the source and target data distributions $p_{\mathrm{s}}$ and $p_{\mathrm{t}}$ using a discrepancy measure between probability distributions, for example an integral probability metric (IPM) or a $\phi$-divergence Sriperumbudur et al. (2009). While $\phi$-divergences like the Kullback-Leibler divergence are well-known, they are often hard to compute for high-dimensional distributions. Wu et al. (2020); Nguyen et al. (2021) suggest that these kinds of measures will correlate with transfer performance, as measured by generalization error on the target task. However, we argue that a meaningful measure of transfer performance must compare to a scratch trained baseline, not target task performance alone. IPMs, such as the Wasserstein-1 Distance, Dudley Metric, and Kernel Mean Discrepency, are bona fide metrics on the space of probability distributions and are of theoretical importance in optimal transport, statistics, and probability theory. Using ideas from optimal transport, (Alvarez-Melis & Fusi, 2020) attempt to correlate transfer performance with the Wasserstein distance. While it may seem that closeness of task distributions should correlate with transfer performance, we show that this is not necessarily the case. In particular, we select a member of each family and prove that, within our model, one can achieve positive transfer ($\mathcal{T}>0$) with distributions that are arbitrarily far apart.
Two functions representable with the same features can be “far apart”. We formalize this notion with the following theorem.
Assumption 2.1.
Theorem 2.2.
Assume 2.1. Then for any $f\in\Phi$, and any $\delta>0$ there exists $g\in\Phi$ such that
$\gamma_{\beta}(p_{f},p_{g})\geq\delta$
$D_{\mathrm{KL}}(p_{f}\|p_{g})\geq\delta$
where $D_{\mathrm{KL}}(p_{f}\|p_{g})$ is the Kullback Leibler divergence.
We prove this theorem in Appendix B.1. We note that this relationship theorem also holds for any IPM over a function class that is larger than the class of Bounded Lipschitz functions. In particular, the theorem holds for the Monge-Kantorovich ($W_{1}$) metric, since any function that satisfies $\lVert f\rVert_{\mathrm{BL}}\leq 1$ also satisfies $\lVert f\rVert\leq_{L}1$. Because it is simpler to compute $W_{1}$ numerically, in Fig. 5 we plot the Wasserstein distance, not the Dudley metric.
Theorem 2.2 demonstrates that for a given source distribution, one can always find a target distribution generated from the same feature space that is arbitrarily distant with respect to these metrics. However, despite such distance, intuitively, if the source and target functions lie in the same feature space and pretraining creates a basis for this space, transfer to the target task should likely be positive. We show this is indeed the case for deep linear networks in the following section.
@@ 3 Deep linear networks: an exactly solvable model
As an analytically solvable model of transfer learning we consider a deep linear network with $L$ layers
@@
where $\bm{W}_{l}\in\mathbb{R}^{d_{l-1}\times d_{l}}$ for $l\in[1,2,\dots L-1]$ and $\bm{W}_{L}\in\mathbb{R}^{d_{L-1}\times 1}$. For notational convenience we have renamed $\bm{c}$ in (3) as $\bm{W}_{L}$ and for simplicity we set $d_{0}=d_{1}=\dots=d_{L-1}=d$, the dimension of the data. The parameter matrices are initialized as $\bm{W}_{l}(0)=\alpha\bar{\bm{W}}_{l}$ where $\alpha\in\mathbb{R}$. Since transfer learning relies on learning features in the source task, we initialize the network in the feature learning regime $\alpha\to 0$. In the following, we assume:
@@
We assume that the source and target functions are each linear functions in $L_{2}(\mathbb{R}^{d},p)$; equivalently, $f_{\mathrm{s}}^{*}(\bm{x})=\bm{\beta}_{\mathrm{s}}^{T}\bm{x}$, $f_{\mathrm{t}}^{*}(\bm{x})=\bm{\beta}_{\mathrm{t}}^{T}\bm{x}$ with $\lVert\bm{\beta}_{\mathrm{s}}\rVert_{2}^{2}=\lVert\bm{\beta}_{\mathrm{t}}\rVert_{2}^{2}$.
@@ 3.1 Pretrained models represent source features
We prove Lemma 3.3 in Appendix B.2. While this result establishes recovery of the ground truth on the source task, it does not describe the feature space of the pretrained model, which is relevant for transferability. To this end, following Yun et al. (2021), we show that in the feature learning regime $\alpha\to 0$, the hidden features of the model sparsify to those present in the source task.
@@
We prove Theorem 3.4 in Appendix B.3. Theorem 3.4 demonstrates that after pretraining in the feature learning regime, the $d$-dimensional feature space of the model parsimoniously represents the ground truth function in a single, rank-one component. We refer to this phenomenon as feature sparsification, which is a hallmark of the feature learning regime, and has important consequences for transferability, particularly in the linear setting Section 3.3.
@@ 3.2 Scratch trained models represent minimum norm solutions
For the empirical training objective 5, there are multiple zero training error solutions when the model is overparameterized $\gamma<1$. As noted in Yun et al. (2021) and Atanasov et al. (2021), there is an implicit bias of gradient flow to the minimum norm solution when $\alpha\to 0$
@@
We prove Theorem 3.5 in Appendix B.4. Knowing the final predictor of the empirical training also allows us to compute the generalization error of the scratch trained model
@@
Theorem (3.6) is a known result for linear regression (Hastie et al., 2022; Canatar et al., 2021; Belkin et al., 2019; Advani & Ganguli, 2016; Mel & Ganguli, 2021; Bartlett et al., 2020), but we provide a proof based on random projections and random matrix theory in Section B.5. This expression exhibits double descent behavior: in the overparameterized regime, the generalization error first decreases, then becomes infinite as $\gamma\to 1$, while in the underparameterized regime, the generalization error monotonically decreases with increasing $\gamma$. As we will see in Section 3.3, this double descent behavior leads to two distinct regions in the transferability phase diagram.

### v2

@@ Abstract
With the emergence of large-scale pre-trained neural networks, methods to adapt such “foundation” models to data-limited downstream tasks have become a necessity. Fine-tuning, preference optimization, and transfer learning have all been successfully employed for these purposes when the target task closely resembles the source task, but a precise theoretical understanding of “task similarity” is still lacking. We adopt a feature-centric viewpoint on transfer learning and establish a number of theoretical results that demonstrate that when the target task is well represented by the feature space of the pre-trained model, transfer learning outperforms training from scratch. We study deep linear networks as a minimal model of transfer learning in which we can analytically characterize the transferability phase diagram as a function of the target dataset size and the feature space overlap. For this model, we establish rigorously that when the feature space overlap between the source and target tasks is sufficiently strong, both linear transfer and fine-tuning improve performance, especially in the low data limit. These results build on an emerging understanding of feature learning dynamics in deep linear networks, and we demonstrate numerically that the rigorous results we derive for the linear case also apply to nonlinear networks.
Machine Learning theory, transfer learning, feature learning, random matrix theory
@@ 1 Introduction
State-of-the-art neural network models have billions to trillions of parameters and are trained on datasets of a similar scale. The benefits of dataset scale are manifest in the astounding generalization capability of these foundation models (Bahri et al., 2024). For many applications, however, datasets of the scale used for natural language processing or computer vision are difficult, if not impossible, to generate. To alleviate the problem of inadequate dataset scale, the representations of a foundation model seem to provide a useful inductive bias for adaptation to a target task. While they are now ubiquitous, transfer learning methods lack a solid theoretical foundation or algorithmic design principles. As such, it remains difficult to predict when—and with which approach—transfer learning will outperform training on the target task alone. By studying an analytically solvable model, we find that the feature space learned during pretraining is the relevant object for predicting transfer performance. Of course, adopting a feature-centric viewpoint creates model-specific challenges because unambiguously identifying learned features remains an outstanding and difficult characterization problem for deep neural networks. For this reason, in this work we focus on deep linear networks trained with gradient flow, as feature learning dynamics are well-understood in this setting [^1] All code to reproduce the results in this paper can be found at https://github.com/javantahir/features_are_fate. We develop an intuitive understanding of linear transfer and full fine-tuning in this model. In contrast to other recent work, we quantify transfer performance relative to training on the target task alone and precisely identify when transfer learning leads to improved performance, effectively building a phase diagram for transfer efficiency. Finally, in numerical experiments, we show that this picture holds qualitatively for nonlinear networks as well.
Related Work
Theoretical aspects of transfer learning
@@
A number of recent works have studied theoretical aspects of transfer learning, focusing on the risk associated with various transfer algorithms. Wu et al. (2020) use information theory to derive bounds on the risk of transfer learning using a mixture of source and target data. Shilton et al. (2017) analyze transfer in the context of Gaussian process regression. Tripuraneni et al. (2020) work in a fairly general setting, and derive bounds on the generalization error of transferred models through a complexity argument, highlighting the importance of feature diversity among tasks. Aminian et al. (2024) study transfer learning in highly overparameterized models, including one hidden layer neural networks, and derive bounds on the excess risk. Bu et al. (2021) study the excess risk of transferred models optimized with the Gibbs algorithm and highlight a bias-variance interpretation of the generalization performance. Liu et al. (2019); Neyshabur et al. (2020) study transfer learning from the perspective of the loss landscape and find that transferred models often find flatter minima than those trained from scratch. Consistent with our feature-centric viewpoint, Kumar et al. (2022) show that fine-tuning can distort the pretrained features, leading to poor out of distribution behavior. Integral probability metrics on the dataset and target have been used to predict transfer learning performance. For example, Wu et al. (2020); Nguyen et al. (2021); Alvarez-Melis & Fusi (2020) suggest that these metrics correlate with generalization error on the target task. However, we prove in Appendix A, that integral probability metrics such as the Wasserstein-1 Distance, Dudley Metric and $\phi-$divergences such as the KL-divergence need not correlate with transfer efficiency, emphasizing that transferrability is a property of the learned feature space and not the source and target datasets alone.
Transfer learning in solvable models
@@
Similar to our approach, several theory works have worked with analytically tractable models to more precisely characterize transfer performance. Lampinen & Ganguli (2018); Atanasov et al. (2021); Shachaf et al. (2021) also study transfer learning in deep linear networks, but focus on the generalization error alone, not the transferability relative to a scratch trained baseline, which obfuscates the conditions for transfer learning to be beneficial. Gerace et al. (2022) studies transfer learning with small nonlinear networks with data generated from a “hidden manifold“ (Goldt et al., 2020) and find it to be effective when tasks are very similar, and data is scarce, but do not theoretically describe regions of negative transfer. Saglietti & Zdeborova (2022) studies knowledge distillation in a solvable model, which can be viewed as a special case of transfer learning. Ingrosso et al. (2024) study transfer learning in a model similar to ours using the replica method and similarly conclude that a feature-based metric for task similarity is predictive of transfer performance.
Feature learning
The feature space of a neural network evolves in the feature learning regime, which should be contrasted with the static neural tangent kernel limit, in which networks regress functions onto a set of basis functions which are fixed at initialization (Jacot et al., 2018). Recently, there has been an explosion of interest in understanding dynamics in these two regimes of neural network optimization. Woodworth et al. (2020); Atanasov et al. (2021); Yang & Hu (2021); Kunin et al. (2024); Yun et al. (2021); Chizat (2020) focus on the role of initialization, learning rate, and implicit bias in feature learning. Petrini et al. (2022) highlights the potential for overfitting when training in the feature learning regime. Allen-Zhu & Li (2020a, b) give a fairly complete treatment of feature learning for synthetic datasets whose features are readily identifiable. Chen et al. (2023) study a synthetic dataset closely related to that in Allen-Zhu & Li (2020a) and describe how generalizable features are important to out of distribution generalization, a conclusion we also draw in the context of transfer learning.
Our contributions
  • Since transfer performance depends on the learned representation of source task, we study deep linear networks trained on a linear target function, a setting in which training dynamics, implicit bias, and generalization error can be described rigorously.
  • Within this model, we analytically compute a phase diagram that illustrates how transfer learning performs relative to training from scratch on a given task.
  • We prove that simple diagnostics, such as distributional measures of source-target distance are insufficient for predicting the success of transfer learning and advance the idea that task similarity should be measured in the space of features instead.
  • We also compute the transfer phase diagram for nonlinear neural networks and show that the same picture applies to the reproducing kernel Hilbert space (RKHS) associated with the nonlinear features of the pre-trained network.
@@ 2 General theoretical setting
We begin by introducing the general theoretical framework under which we study transfer learning. We consider source and target regression tasks defined by probability distributions $p_{\mathrm{s}}(\bm{x},y)$ and $p_{\mathrm{t}}(\bm{x},y)$ over inputs $\bm{x}\in\mathbb{R}^{d}$ and labels $y\in\mathbb{R}$. We focus on concept drift, in which $p_{\mathrm{s}}(\bm{x},y)=p(\bm{x})p_{\mathrm{s}}(y\mid\bm{x})$ and $p_{\mathrm{t}}(\bm{x},y)=p(\bm{x})p_{\mathrm{t}}(y\mid\bm{x})$ for the same input distribution $p(\bm{x})$. The labels are generated from noisy samples of source and target functions $y_{\textrm{s}}=f_{\textrm{s}}^{*}(\bm{x})+\epsilon_{\rm s}\textrm{ and }y_{\textrm{t}}=f_{\textrm{t}}^{*}(\bm{x})+\epsilon_{\rm t}$ where $f_{\textrm{s}}^{*}(\bm{x}),f_{\textrm{t}}^{*}(\bm{x})\in L_{2}(p(\bm{x}))$ and $\epsilon_{\rm s},\epsilon_{\rm t}\sim\mathcal{N}(0,\sigma^{2})$. During pretraining, we train a model with parameters $\bm{\Theta}=(\bm{c},\bm{\theta})$ of the form
@@
where $\hat{\mathbb{E}}_{p}(h(\bm{x},y))=\frac{1}{n}\sum_{i=1}^{n}h(\bm{x}_{i},y_{i})$ is the expectation over the empirical distribution of $\mathcal{D}$. We focus on two widely employed transfer methods, linear transfer and fine-tuning. In linear transfer, the pretrained features $\phi(\bm{x},\bm{\theta})$ are frozen and only the output weights $\bm{c}$ are trained on the target task. In fine-tuning, the entire set of parameters $\bm{\Theta}$ are trained from the pretrained initialization on the target task. In Appendix C, we explore the case when the source dataset is also finite, and show that for fine-tuning, the qualitative picture discussed below is unchanged. To optimize the loss functions (2) and (3), we use gradient flow,
@@
where we have set the learning rate equal to unity for the purpose of analysis. To assess the performance of transfer learning we compare the generalization of the transferred model to a scratch-trained model with the same architecture (1) trained on the target same data from a random initialization. We introduce the transferability to quantify this relationship:
@@
$\mathcal{R}=\mathbb{E}_{p(\bm{x})}[((f(\bm{x},\bm{\Theta})-f^{*}(\bm{x}))^{2}].$ (6)
@@ 3 Deep linear networks: an exactly solvable model
Because insight into feature learning is required to understand the dynamics of transfer learning, we first consider an analytically solvable model of transfer learning using deep linear networks. This model gives us direct access to the evolution of the time-dependent feature space and its distortion under various transfer learning schemes. Let $f$ denote a deep linear network with $L$ layers
@@
where $\bm{W}_{l}\in\mathbb{R}^{d_{l-1}\times d_{l}}$ for $l\in[1,2,\dots L-1]$ and $\bm{W}_{L}\in\mathbb{R}^{d_{L-1}\times 1}$. For notational convenience we have renamed $\bm{c}$ in (1) as $\bm{W}_{L}$ and for simplicity we set $d_{0}=d_{1}=\dots=d_{L-1}=d$, the dimension of the data. The parameter matrices are initialized as $\bm{W}_{l}(0)=\alpha\bar{\bm{W}}_{l}$ where $\alpha\in\mathbb{R}$. The matrices $\bm{W}_{l}(0)$ additionally satisfy (19), which is a technical assumption that generalizes common initialization schemes such as He initialization (Yun et al., 2021; He et al., 2015). Since transfer learning relies on learning features in the source task, we initialize the network in the feature learning regime $\alpha\to 0$. In the following, we assume:
@@
We assume that the source and target functions are each linear functions in $L_{2}(\mathbb{R}^{d},p)$; equivalently, $f_{\mathrm{s}}^{*}(\bm{x})=\bm{\beta}_{\mathrm{s}}^{T}\bm{x}$, $f_{\mathrm{t}}^{*}(\bm{x})=\bm{\beta}_{\mathrm{t}}^{T}\bm{x}$ with $\lVert\bm{\beta}_{\mathrm{s}}\rVert_{2}^{2}=\lVert\bm{\beta}_{\mathrm{t}}\rVert_{2}^{2}=1$.
@@
Linear networks have the advantage of analytic tractability, but we note that the representation capacity of these models is limited to affine transformations. Furthermore, the expressiveness of the model is independent of the number of layers. As a result, this model may fail to capture aspects of transfer learning that depend strongly on depth separation  (Telgarsky, 2016; Daniely, 2017) or other nonlinear phenomena. However, overparameterized linear models, recapitulate many phenomena observed in deep learning, including double descent (Nakkiran et al., 2021; Belkin et al., 2019), scaling laws (Bahri et al., 2024), feature learning (Vyas et al., 2024; Atanasov et al., 2021) and, as we show, the impact of feature learning on transfer efficiency.
@@ 3.1 Pretrained models represent source features
We prove Lemma 3.3 in Appendix D.2. While this result establishes recovery of the ground truth on the source task, it does not describe the feature space of the pretrained model, which is relevant for transferability. To this end, following Yun et al. (2021), we show that in the feature learning regime $\alpha\to 0$, the hidden features of the model sparsify to those present in the source task.
@@
We prove Theorem 3.4 in Appendix D.3. Theorem 3.4 demonstrates that after pretraining in the feature learning regime, the $d$-dimensional feature space of the model parsimoniously represents the ground truth function in a single, rank-one component. We refer to this phenomenon as feature sparsification, which is a hallmark of the feature learning regime, and has important consequences for transferability, particularly in the linear setting Section 3.3.
@@ 3.2 Scratch trained models represent minimum norm solutions
For the empirical training objective (3), there are multiple zero training error solutions when the model is overparameterized $\gamma<1$. As noted in Yun et al. (2021) and Atanasov et al. (2021), there is an implicit bias of gradient flow to the minimum norm solution when $\alpha\to 0$
@@
We prove Theorem 3.5 in Appendix D.4. Knowing the final predictor of the empirical training also allows us to compute the generalization error of the scratch trained model
@@
Theorem (3.6) is a known result for linear regression (Hastie et al., 2022; Canatar et al., 2021; Belkin et al., 2019; Advani & Ganguli, 2016; Mel & Ganguli, 2021; Bartlett et al., 2020), but we provide a proof based on random projections and random matrix theory in Section D.5. This expression exhibits double descent behavior: in the overparameterized regime, the generalization error first decreases, then becomes infinite as $\gamma\to 1$, while in the underparameterized regime, the generalization error monotonically decreases with increasing $\gamma$. As we will see in Section 3.3, this double descent behavior leads to two distinct regions in the transferability phase diagram. The fact that scratch-trained performance can be arbitrarily bad is a result of the implicit regularization of gradient flow on this model. This effect can be eliminated by appropriately regularizing the scratch-trained model with weight decay. In the interest of analytic tractability we do not include regularization when training from scratch, but we explore its effects in simulation in Appendix G Fig. 5